# Cooling and heating regions of Joule-Thomson expansion for AdS black holes: Maxwell-power-Yang-Mills and Kerr Sen black holes

### v1

@@ Abstract
In this paper, we explore the Joule-Thomson expansion (JTE) process for the Einstein-Power-Young-Mills (EPYM) and the AdS Kerr Sen (AKS) black holes. We study the effect of free parameters on the Joule-Thomson coefficient (JTC), the inversion curve, and the $T_{i}^{min}/T_{c}$. The isenthalpic curves of the AKS black hole show cooling or heating behavior depending on the inversion curve, which is affected by the mass and the parameters $b$ and $a$ of the black hole. If we assume the parameter $b$ to be zero, the results reduce to the Kerr–AdS black holes[1]. In [2, 3], for the Einstein–Power–Yang–Mills AdS black hole with $q>1$ and $n=2$, the $T_{i}^{min}/T_{c}$ is $1/2$. But in this paper, for the AdS-Maxwell-power-Yang-Mills black hole, when $q>1$, the $T_{i}^{min}/T_{c}$ is almost equal to $1/2$ for the increase of Maxwell’s charge $C$, and when $q=1/2$, the $T_{i}^{min}/T_{c}$ is equal to $1/2$ for all values of $C$. Also, when $11/2<q<$ 1, the $T_{i}^{min}/T_{c}$ is close to the value of $1/2$, and finally when $20<q<1/$ 2, the values of $T_{i}^{min}/T_{c}$ move away from the value of $1/2$, that is, they become smaller. For the AKS black hole, we found that for free parameters $a=0.00951$ and $b=0.00475$, the $T_{i}^{min}/T_{c}$ is almost $1/2$. Finally, we compare our findings with others in the literature and summarize our results in Tables 1-5.
Keywords: JTE, AdS-Maxwell-power-Yang-Mills black holes, AdS Kerr Sen black holes
@@ 1 Introduction
The insights provided by astronomy, astrophysics, and experimental cosmology suggest that cosmic structures adhere to the same principles and foundations observed on Earth, albeit with minor variations and necessary simplifications. This structural and fractal similarity is a crucial tool for human ideation and cognition, enabling us to decode more aspects of this global pattern daily. Researchers leverage this understanding in hypothesizing theoretical models, which may not yet be empirically verifiable. They construct various models based on mathematical logic and physical laws, exploring all potential scenarios to predict the most plausible ideas consistent with the formation of the enigmatic, infinite universe. The existing black hole models, are actually one of the most obvious examples of this form of inference. Despite being one of the most elusive cosmic entities, they are studied based on this pattern. The comparison of a black hole’s gravitational behavior with a thermodynamic ensemble in the 1970s gave rise to a significant branch of black hole physics, namely black hole thermodynamics. For instance, the four laws of black hole mechanics bear a striking resemblance to the laws of thermodynamics [4, 5]. Similarly, a black hole’s surface area is analogous to entropy in thermodynamics, and its surface gravity is comparable to temperature[4, 5]. The introduction of Maldacena duality, also known as the AdS/CFT correspondence, established a deeper connection between black hole thermodynamics and this duality, offering profound insights into quantum gravity. In the context of AdS/CFT, a black hole’s entropy in the bulk AdS space correlates with the entropy of the corresponding CFT on the boundary, known as the Bekenstein-Hawking entropy [6]. Moreover, the black hole’s temperature is related to the CFT’s temperature[6, 7, 8, 9].
@@
This correspondence provides a powerful tool to study the quantum aspects of gravity and black holes using the methods of quantum field theory. In AdS space, there is a Hawking-Page phase transition between a stable large black hole and a thermal gas[10]. This phase transition is a first-order transition that occurs when the temperature of the system reaches a critical value, where the free energy of the black hole becomes lower than that of the thermal gas[10]. This phase transition can be interpreted as a confinement deconfinement phase transition of a gauge field[11]. The Hawking-Page phase transition can be seen as a transition from a deconfined phase in the thermal gas to a confined phase in the black hole[11]. When the AdS black holes have electric charge, they exhibit rich phase structures that were studied by Chamblin et al[12, 13]. They found that the phase transition behavior of charged AdS black holes resembles the liquid-gas phase transition in a van der Waals system[14]. In the extended phase space where the cosmological constant is treated as pressure[15], the P-V critical behavior of charged AdS black holes was investigated and it was shown that they have a similar analogy to the van der Waals liquid-gas system. In addition to the phase transition and critical phenomena[16, 17, 18, 19, 20, 21, 22, 23], the analogy between the black holes and the van der Waals system was also creatively applied to the well-known JTE process[24] recently. This means that the JTC can be used to study the thermodynamics of black holes as well. For example, the isenthalpic expansion process is the analogue of the JTE process for black holes, where the black hole mass is constant while the black hole pressure and volume are changed. The black hole pressure is related to the cosmological constant, and the black hole volume is related to the horizon radius.In this case, the inversion curve is the curve that separates the regions where the black hole temperature increases or decreases as the pressure decreases.One of the intriguing features of the inversion curves for black hole systems is that they have only positive slopes, unlike the van der Waals system, which has both positive and negative slopes. For example, for charged AdS black holes [24] and Kerr-AdS black holes [1], the isenthalpic expansion process and the inversion curve have been analyzed and observed that the inversion curve was found to have a positive slope, meaning that the black hole temperature always decreases as the pressure decreases. Then the analysis was generalized to other types of AdS black holes, such as quintessence charged AdS black holes[25], holographic superfluids [26], charged AdS black holes in f(R) gravity [27], AdS black holes with a global monopole [28], and AdS black holes in Lovelock gravity [29]. For further study, you can see also [30, 31, 32, 33, 34, 35, 36, 37, 2, 38].
@@
All the results showed that the inversion curves for all these black hole systems have only positive slopes. We are interested in exploring whether this feature is universal for all black hole systems, or whether there are other effects that can alter it. To this end, we will focus on two types of AdS black holes: AdS-Maxwell-power-Yang-Mills and AdS Kerr Sen. These black holes have additional parameters that can affect their thermodynamic behavior and phase transitions. AdS-Maxwell-power-Yang-Mills black holes are black holes that have a non-linear electromagnetic field, which is described by a power-law function of the field strength[39]. This field can be seen as a generalization of the Maxwell field, which is the standard model of electromagnetism. AdS Kerr Sen black holes are black holes that have electric charge and angular momentum in a low-energy limit of heterotic string theory[40]. This theory is a type of string theory that combines the features of bosonic and supersymmetric strings. These black holes have different properties and characteristics than the standard AdS black holes, such as the existence of a dilaton field, which is a scalar field that couples to the electromagnetic field and the curvature. We will investigate how these parameters affect the inversion curves and the JTE process for these black holes, and compare them with the previous results for other black hole systems.
@@
The structure of this paper is as follows. In Sec.II, we give a brief overview of the JTE. In Sec.III and Sec.IV, We introduce briefly AdS-Maxwell-Power-Yang-Mills and the AdS Kerr Sen black holes and their thermodynamic properties.In Sec.V, we study the JTE process for these black holes, derive an explicit expression for the JTC, and analyze and discuss the effect of the parameters of each model on the inversion curves. In Sec.VI, we present our conclusion and discussion.
@@ 2 Joule-Thomson expansion
In classical thermodynamics, a throttling process or the JTE process, discovered in 1852, is a method of cooling or heating a system by changing its pressure and volume, without adding or removing heat. In this process, the high-pressure gas passes through a porous plug into a region with a low pressure, while keeping the enthalpy constant. Since it is a constant-enthalpy process, it can be used to experimentally measure the lines of constant enthalpy (isenthalps) on the (p, T) diagram of a system. Combined with the specific heat capacity at constant pressure, it allows the complete measurement of the thermodynamic potential for the gas[41].
@@
In this method, The main goal is to investigate the behavior of the coefficient that describes the temperature change during the expansion or compression of a system at constant enthalpy, which is denoted by $\mu$ and is known as the JTC,
$\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{H}=\frac{1}{C_{P}}\big{[}T(\frac{\partial V}{\partial T}\big{)}_{P}-V\big{]}.$
@@
If the above coefficient is positive, as a result of pressure reduction, the temperature will decrease. In other words, the expansion of the gas causes cooling and the compression of the gas under investigation causes it to heat up. In other words, the positive JTC indicates the same direction of temperature and pressure. Whereas, if the JTC is negative, a decrease in pressure causes an increase in temperature.
@@
The JT inversion temperature,which determined by setting $\mu=0$, is the temperature at which the sign of the JTC changes. Most real gases have an inversion point. The temperature of this point depends on the gas pressure before expansion.
$T_{i}=V\big{(}\frac{\partial T}{\partial V}\big{)}_{P}.$
@@
The important point is that if you plot the JTC on the (p,T) diagram, a closed parabolic curve is created. In simpler terms, the inversion temperature is placed on the boundary of the curve of temperature changes in terms of pressure. At this temperature, the JTC changes from negative to positive. At a given pressure, the isopressure-temperature line intersects the drawn curve at two different points. These two points are called low temperature and high temperature of inversion. In fact, at temperatures higher and lower than these two temperatures, the sign of the JTC is negative and between these two temperatures, the sign of the JTC is positive. According to the above, the interesting phenomenon in this process is that the (p, T) diagram has two regions: one where the gas cools down and one where the gas heats up. These regions are separated by the inversion curve,the curve that shows the points where the system temperature does not change during the expansion process, that divides the graph into two regions: the cooling region and the heating region. The cooling region is where the gas temperature decreases as the pressure decreases, and the heating region is where the gas temperature increases as the pressure decreases. The inversion curve depends on the type of the gas and its initial conditions. If with respect to zero coefficient for ideal gases, we choose the van der Waals system, which is a more realistic model than the ideal gas, and takes into account the finite size and the attractive forces of the molecules, we find that for the van der Waals system, the inversion curves have both positive and negative slopes, forming a circle in the pressure axis. The inversion curve for the van der Waals system has a negative slope in the low-pressure region, where the attractive forces dominate, and a positive slope in the high-pressure region, where the repulsive forces dominate.
@@
3 Case I: AdS-Maxwell-power-Yang-Mills black holes
The AMPYM black holes are rooted in the study of black hole solutions in the context of supergravity theories, especially in anti-de Sitter (AdS) space. The study of black holes in AdS space is important for various reasons, including its relevance to string theory and the AdS/CFT correspondence, which is a duality between gravitational theories in AdS space and field theories defined on its boundary. These black holes are solutions to the equations of motion of supergravity theories with additional matter fields, such as Maxwell fields and power-Yang-Mills fields, within the context of AdS space that arises from a generalization of the Einstein-Maxwell theory with a negative cosmological constant and a non-Abelian gauge field.The history of AMPYM black holes can be traced back to the discovery of the first black holes in Einstein-Yang-Mills theory, which were considered in the works of Yasskin [42] and Kasuya [43]. It should be noted that in the non-Abelian case there are various gauge groups, but to obtain black hole solutions it is necessary to choose a specific form for the gauge potential. One of the simplest forms that nevertheless allowed to derive interesting and important results is the so-called Wu-Yang ansatz, which leads to magnetic-type solutions. [44, 45, 46, 47]. Of course, an interesting point that can be mentioned is that the primary black holes with non-Abelian fields, which were studied in the late 80s [48, 49, 50, 51], were found to be unstable in the case of asymptotically flat geometry [52, 53], but later black hole solutions were obtained in the AdS case, which were shown to be stable [54, 55, 56, 57]. The AMPYM black holes have been extensively studied in the context of theoretical physics, particularly in the context of holography and its applications to understanding strongly coupled field theories. In recent years, the study of AMPYM black holes has continued to be an active area of research, with a focus on understanding their thermodynamic properties, phase transitions, and connections to gauge/gravity duality. These black holes have been studied as models for strongly coupled systems in the dual field theories, providing valuable insights into nonperturbative phenomena in quantum field theories. Researchers have investigated various aspects of AMPYM black holes, such as their stability, entropy, and critical behavior. Their thermodynamic properties have been of particular interest, as they exhibit rich phase structures and can undergo phase transitions similar to those observed in condensed matter systems. Furthermore, the holographic interpretation of these black holes has led to fruitful connections between gravitational physics and field theory, shedding light on fundamental questions in quantum gravity and strongly coupled systems[58, 59, 60]. This section provides a summary of the thermodynamics of N-dimensional Einstein-Maxwell-power-Yang-Mills gravity with a cosmological constant $\Lambda$. This theory of gravity is based on the following action, which we will explain in more details. So we will have,
@@ 3 Case I: AdS-Maxwell-power-Yang-Mills black holes
The trace element is represented by $Tr(.)=\Sigma_{a=1}^{(N-1)(N-2)/2}(.)$, with $R$ as the Ricci scalar and $q$ as a real positive parameter. The Yang–Mills and Maxwell fields are defined accordingly [4],
$\begin{split}F_{\mu\nu}^{(a)}=\partial_{\mu}A_{\nu}^{(a)}-\partial_{\nu}A_{\mu}^{(a)}+\frac{1}{2\sigma}C^{(a)}_{(b)(c)}A^{b}_{\mu}A^{c}_{\nu},\end{split}$ (2)
$\begin{split}F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},\end{split}$ (3)
@@
$C^{(a)}_{(b)(c)}$ represents the structure constants of the $(N-1)(N-1)/2$ parameter Lie group $G$, while $\sigma$ denotes the coupling constant.$A^{(a)}_{\mu}$ refers to the $SO(N-1)$ gauge group Yang-Mills potentials, with $A_{\mu}$ representing the conventional Maxwell potential [39]. The metric solution corresponding to the $N$-dimensional spherically symmetric line element is as follows [40],
$\begin{split}ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\Omega_{n}^{2}.\end{split}$ (4)
The $d\Omega_{n}^{2}$ denotes the volume of the unit $n$-sphere. In this study, we will direct our attention to the Einstein-Maxwell-power-Yang-Mills theory ($EMPYM$) with $N(=n+2)\geq 4$ and $q\neq(n+1)/4$. The solution to the $N$-dimensional $EMPYM$ black hole with a negative cosmological constant under the condition of $q\neq\frac{n+1}{4}$ is provided [39],
$\begin{split}&f(r)=1-\frac{2m}{r^{n-1}}-\frac{\Lambda r^{2}}{3}+\frac{2(n-1)C^{2}}{nr^{2(n-1)}}+\frac{Q}{r^{4q-2}}\\ &Q=\frac{[n(n-1)Q_{1}^{2}]^{q}}{n(4q-n-1)}.\end{split}$ (5)
@@
It is important to note that the parameter $m$ represents the mass of the black hole, while $C$ and $Q_{1}$ denote the charges of the Maxwell field and Yang-Mills field, respectively. In the extended phase space, the cosmological constant is considered as a thermodynamic pressure $P=-\frac{\Lambda}{8\pi}$, in this case, the Hawking temperature, mass and entropy of the black hole are obtained as follows,
$\begin{split}T=-\frac{C^{2}(n-1)^{2}}{2\pi nr_{+}^{2n-1}}+\frac{2}{3}(n+1)Pr_{+}-\frac{Q(-n+4q-1)}{4\pi r_{+}^{4q-1}}+\frac{n-1}{4\pi r_{+}},\end{split}$ (6)
$\begin{split}M=\frac{n\omega_{n}}{48\pi}\bigg{(}\frac{6C^{2}(n-1)}{nr_{+}^{n-1}}+8\pi Pr_{+}^{n+1}+3Qr_{+}^{n-4q+1}+3r_{+}^{n-1}\bigg{)},\end{split}$ (7)
@@
$\begin{split}S=\frac{\omega_{n}r_{+}^{n}}{4},\qquad\omega_{n}=\frac{2\pi^{\frac{n+1}{2}}}{\Gamma\left(\frac{n+1}{2}\right)},\end{split}$ (8)
where $r_{+}$ and $\omega_{n}$ are the horizon radius and the volume of the unit $n$-sphere respectively.

### v2

# Cooling and heating regions of Joule-Thomson expansion for AdS black holes: Einstein-Maxwell-Power-Yang-Mills and Kerr Sen black holes

@@ Abstract
In this paper, we study the Joule-Thomson Expansion (JTE) process for two types of black holes: AdS-Einstein–Maxwell-Power-Yang–Mills (AEMPYM) and AdS-Kerr-Sen (AKS). Our study focuses on understanding how various parameters influence the Joule-Thomson Coefficient (JTC), the inversion curve, and the ratio of minimum inversion temperature to critical temperature. For the AKS black hole, we observe that the isenthalpic curves can exhibit either cooling or heating behavior. This behavior is determined by the inversion curve, which is affected by the black hole’s mass and specific parameters such as $b$ (parameter signifies the ionic charge of the black hole) and $a$ (rotation parameter). In the case of the AEMPYM black hole, our findings reveal that the ratio of minimum inversion temperature to critical temperature approaches a specific value as Maxwell’s charge increases. This ratio remains constant for certain parameter values, while it varies for others. Specifically, when the parameter $q$ (real positive parameter of AEMPYM black hole) is greater than 1, the ratio is almost equal to 1/2 as Maxwell’s charge (C) increases. When q equals 1/2, the ratio is exactly 1/2 for all values of (C). For values of (q) between 1/2 and 1, the ratio is close to 1/2, and for values of (q) between 0 and 1/2, the ratio decreases, moving away from 1/2. For the AKS black hole, we find that specific parameter values, such as (a = 0.00951) and (b = 0.00475 ), yield a ratio of minimum inversion temperature to a critical temperature that is approximately 1/2. This consistency across different parameter values highlights the robustness of our findings. Finally, we compare our results with those reported in the existing literature, providing a comprehensive summary in detailed tables. This comparison not only validates our findings but also situates them within the broader context of black hole thermodynamics and the Joule-Thomson effect.
Keywords; Joule-Thomson expansion, AdS-Maxwell-power-Yang-Mills black holes, AdS-Kerr-Sen black holes
Corresponding author; Saeed Noori Gashti
@@ 1 Introduction
The insights provided by astronomy, astrophysics, and experimental cosmology suggest that cosmic structures adhere to the same principles and foundations observed on Earth, albeit with minor variations and necessary simplifications. This structural and fractal similarity is a crucial tool for human creativity and cognition, enabling us to decode more aspects of this global pattern daily. Researchers leverage this understanding in hypothesizing theoretical models, which may not yet be empirically verifiable[1]. They construct various models based on mathematical logic and physical laws, exploring all potential scenarios to predict the most plausible ideas consistent with the formation of the mysterious, infinite universe. The existing black hole models are one of the most obvious examples of this form of inference[2]. Despite being one of the most important cosmic entities, they are studied based on this pattern. The comparison of a black hole’s gravitational behavior with a thermodynamic ensemble in the 1970s gave rise to a significant branch of black hole physics, namely black hole thermodynamics. For instance, the four laws of black hole mechanics bear a striking resemblance to the laws of thermodynamics [3, 4]. Similarly, a black hole’s surface area is analogous to entropy in thermodynamics, and its surface gravity is comparable to temperature[3, 4]. The introduction of Maldacena duality, also known as the AdS/CFT correspondence, established a deeper connection between black hole thermodynamics and this duality, offering profound insights into quantum gravity. In the context of AdS/CFT, a black hole’s entropy in the bulk AdS space correlates with the entropy of the corresponding CFT on the boundary, known as the Bekenstein-Hawking entropy [5]. Moreover, the black hole’s temperature is related to the CFT’s temperature[5, 6, 7, 8].
@@
This correspondence provides a powerful tool to study the quantum aspects of gravity and black holes using the methods of quantum field theory. In AdS space, there is a Hawking-Page phase transition between a stable large black hole and a thermal gas[9]. This phase transition is a first-order transition that occurs when the temperature of the system reaches a critical value, where the free energy of the black hole becomes lower than that of the thermal gas[9]. This phase transition can be interpreted as a confinement/deconfinement phase transition of a gauge field[10]. The Hawking-Page phase transition can be seen as a transition from a deconfined phase in the thermal gas to a confined phase in the black hole[10]. When the AdS black holes have an electric charge, they exhibit rich phase structures that were studied by Chamblin et al[11, 12]. They found that the phase transition behavior of charged AdS black holes resembles the liquid-gas phase transition in a van der Waals system[13]. In the extended phase space where the cosmological constant is treated as pressure[14], the Pressure-Volume (P-V) critical behavior of charged AdS black holes was investigated and it was shown that they have a similar analogy to the van der Waals liquid-gas system. In addition to the phase transition and critical phenomena[15, 16, 17, 18, 19, 20, 21, 22], the analogy between the black holes and the van der Waals system was also creatively applied to the well-known JTE process[23] recently. This means that the JTC can be used to study the thermodynamics of black holes as well[24, 25]. For example, the isenthalpic expansion process is the analog of the JTE process for black holes, where the black hole mass is constant while the black hole pressure and volume are changed. The black hole pressure is related to the cosmological constant, and the black hole volume is related to the horizon radius. In this case, the inversion curve is the curve that separates the regions where the black hole temperature increases or decreases as the pressure decreases. One of the intriguing features of the inversion curves for black hole systems is that they have positive slopes, unlike the van der Waals system, which has both positive and negative slopes. For example, for charged AdS black holes [23] and Kerr-AdS black holes [26], the isenthalpic expansion process and the inversion curve have been analyzed and observed that the inversion curve was found to have a positive slope, meaning that the black hole temperature always decreases as the pressure decreases. Then the analysis was generalized to other types of AdS black holes, such as quintessence charged AdS black holes[29], holographic superfluids [30], charged AdS black holes in f(R) gravity [31], AdS black holes with a global monopole [32], and AdS black holes in Lovelock gravity [33]. For further study, you can see in [34, 35, 36, 37, 38, 39, 40, 41, 27, 28, 42].
In general, the study of the thermodynamics of black holes in different frameworks, such as extended, CFT, and RPS phase space, as well as the study of phase transitions, especially their thermodynamic topology, has recently received much attention. a comprehensive review of these concepts is in the literature [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
@@
All the results showed that the inversion curves for all these black hole systems have only positive slopes. We are interested in exploring whether this feature is universal for all black hole systems, or whether other effects can alter it. To this end, we will focus on two types of black holes: AEMPYM and AKS. These black holes have additional parameters that can affect their thermodynamic behavior and phase transitions. AEMPYM are black holes that have a non-linear electromagnetic field, which is described by a power-law function of the field strength[65]. This field can be seen as a generalization of the Maxwell field, which is the standard model of electromagnetism. AKS black holes are black holes that have electric charge and angular momentum in a low-energy limit of heterotic string theory[66]. This theory is a type of string theory that combines the features of bosonic and supersymmetric strings. These black holes have different properties and characteristics than the standard AdS black holes, such as the existence of a dilaton field, which is a scalar field that couples to the electromagnetic field and the curvature. We will investigate how these parameters affect the inversion curves and the JTE process for these black holes, and compare them with the previous results for other black hole systems.
@@
The structure of this paper is as follows. In Sec. 2, we give a brief overview of the JTE. In Sec. 3 and Sec. 4, we introduce briefly AEMPYM and the AKS black holes and their thermodynamic properties. In Sec. 5, we study the JTE process for these black holes, derive an explicit expression for the JTC, and analyze and discuss the effect of the parameters of each model on the inversion curves. In Sec. 6, we present our conclusion and discussion.
@@ 2 Joule-Thomson expansion
In classical thermodynamics, the JTE process, discovered in 1852, is a method of cooling or heating a system by changing its pressure and volume, without adding or removing heat. In this process, the high-pressure gas passes through a porous plug into a region with a low pressure, while keeping the enthalpy constant. Since it is a constant-enthalpy process, it can be used to experimentally measure the lines of constant enthalpy on the $(P,T)$ figure of a system. Combined with the specific heat capacity at constant pressure, it allows the complete measurement of the thermodynamic potential for the gas[67, 34, 35, 27, 28, 26].
@@
In this method, The main goal is to investigate the behavior of the coefficient that describes the temperature change during the expansion or compression of a system at constant enthalpy, which is denoted by $\mu$ and is known as the JTC[67, 34, 35, 27, 28, 26],
$\mu=\big{(}\frac{\partial T}{\partial P}\big{)}_{H}=\frac{1}{C_{P}}\big{[}T(\frac{\partial V}{\partial T}\big{)}_{P}-V\big{]},$
@@
where T, P, V, $C_{P}$, H are determined by temperature, pressure, volume, specific heat capacity at constant pressure, and enthalpy (the mass of the black hole could be considered as the enthalpy, $M\equiv H$, rather than the internal energy of the gravitational system), respectively. If the above coefficient is positive, as a result of pressure reduction, the temperature will decrease. In other words, the expansion of the gas causes cooling and the compression of the gas under investigation causes it to heat up. In other words, the positive JTC indicates the same direction of temperature and pressure. Whereas, if the JTC is negative, a decrease in pressure causes an increase in temperature.[67, 34, 35, 27, 26]
@@
The JT inversion temperature, which is determined by setting $\mu=0$, is the temperature at which the sign of the JTC changes. Most real gases have an inversion point. The temperature of this point depends on the gas pressure before expansion[67, 34, 35, 27, 26].
$T_{i}=V\big{(}\frac{\partial T}{\partial V}\big{)}_{P},$
@@
where $T_{i}$ is inversion temperature. The important point is that if you plot the JTC on the $(P,T)$ diagram, a closed parabolic curve is created. In simpler terms, the inversion temperature is placed on the boundary of the curve of temperature changes in terms of pressure[67, 34, 35, 27, 26]. At this temperature, the JTC changes from negative to positive. At a given pressure, the line intersects the drawn curve at two different points. These two points are called low temperature and high temperature of inversion. In fact, at temperatures higher and lower than these two temperatures, the sign of the JTC is negative and between these two temperatures, the sign of the JTC is positive[67, 34, 35, 27, 26]. According to the above, the interesting phenomenon in this process is that the (P, T) diagram has two regions: one where the gas cools down and one where the gas heats up[67, 34, 35, 27, 26]. These regions are separated by the inversion curve, the curve that shows the points where the system temperature does not change during the expansion process, which divides the graph into two regions: the cooling region and the heating region. The cooling region is where the gas temperature decreases as the pressure decreases, and the heating region is where the gas temperature increases as the pressure decreases[67, 34, 35, 27, 26]. The inversion curve depends on the type of the gas and its initial conditions. If with respect to zero coefficient for ideal gases, we choose the van der Waals system, which is a more realistic model than the ideal gas, and takes into account the finite size and the attractive forces of the molecules, we find that for the van der Waals system, the inversion curves have both positive and negative slopes, forming a circle in the pressure axis. The inversion curve for the van der Waals system has a negative slope in the low-pressure region, where the attractive forces dominate, and a positive slope in the high-pressure region, where the repulsive forces dominate[67, 34, 35, 27, 26].
@@
3 Case I: AdS-Einstein-Maxwell-Power-Yang-Mills black holes
AEMPYM black holes are studied within supergravity theories, particularly in AdS space, due to their relevance to string theory and the AdS/CFT correspondence. These black holes are solutions to supergravity equations with additional matter fields like Maxwell and power-Yang-Mills fields, extending the Einstein-Maxwell theory with a negative cosmological constant and a non-Abelian gauge field. The history of AEMPYM black holes dates back to the first black holes in Einstein-Yang-Mills theory by Yasskin and Kasuya. Initially, non-Abelian black holes were unstable in asymptotically flat geometries, but later found stability in AdS space. The Wu-Yang ansatz is a key form for deriving magnetic-type solutions. Recent research focuses on their thermodynamic properties, phase transitions, and connections to gauge/gravity duality. These black holes model strongly coupled systems in dual field theories, offering insights into nonperturbative phenomena. Their rich phase structures and holographic interpretations link gravitational physics with field theory, addressing fundamental questions in quantum gravity and strongly coupled systems[68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 65].
Here, we investigate the thermodynamic properties of N-dimensional gravity theories that combine Einstein’s gravity, Maxwell’s electromagnetism, and power-Yang-Mills fields. The presence of a cosmological constant $\Lambda$ further enriches the dynamics of these gravitational systems. The theory is grounded in the following action, which encapsulates the fundamental interactions[3],
@@ 3 Case I: AdS-Einstein-Maxwell-Power-Yang-Mills black holes
In the context of N-dimensional gravity theories, we encounter the trace element and field definitions that play essential roles in understanding spacetime dynamics, $Tr(.)=\sum_{a=1}^{(N-1)(N-2)/2}(.)$ Here, $R$ represents the Ricci scalar, and $q$ is a real positive parameter. The summation involves a combination of indices, capturing the intricate geometric properties of the manifold. Yang-Mills and Maxwell Fields are fundamental in describing the interactions within the gravitational system. The Yang-Mills field corresponds to non-abelian gauge theories, capturing the strong force interactions. The Maxwell field represents electromagnetism, describing the behavior of charged particles. In summary, the trace element and the interplay between Yang-Mills and Maxwell fields contribute to the rich tapestry of N-dimensional gravity. These concepts are crucial for unraveling the underlying physics and exploring the intricate fabric of spacetime. The Yang-Mills and Maxwell fields are defined as [3],
$\begin{split}F_{\mu\nu}^{(a)}=\partial_{\mu}A_{\nu}^{(a)}-\partial_{\nu}A_{\mu}^{(a)}+\frac{1}{2\sigma}C^{(a)}_{(b)(c)}A^{b}_{\mu}A^{c}_{\nu},\end{split}$ (3.2)
$\begin{split}F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},\end{split}$ (3.3)
@@
The structure constants of the Lie group $(G)$, which has $((N-1)(N-1)/2)$ parameters, are denoted by $(C^{(a)}_{(b)(c)})$. The coupling constant is represented by $(\sigma)$, and the Yang-Mills potentials for the $(SO(N-1))$ gauge group are given by $(A^{(a)}_{\mu})$. In this context, $(A_{\mu})$ signifies the conventional Maxwell potential as discussed in reference[65]. The metric solution that corresponds to the line element of an ( N )-dimensional spherically symmetric space is detailed in reference[66].
$\begin{split}ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\Omega_{n}^{2}.\end{split}$ (3.4)
The term $(d\Omega_{n}^{2})$ signifies the volume element on an ( n)-dimensional unit sphere. The scope of this study is the EMPYM theory in dimensions $(N)$, where $(N=n+2)$ and is at least 4, with the constraint that the parameter $(q)$ is distinct from ( $\frac{n+1}{4}$). q is real positive parameter of AEMPYM black hole. A detailed exposition of the ( N )-dimensional EMPYM black hole solution, which includes a negative cosmological constant and conforms to the condition ( $q\neq\frac{n+1}{4}$ ), is presented in the referenced document[65].
$\begin{split}&f(r)=1-\frac{2m}{r^{n-1}}-\frac{\Lambda r^{2}}{3}+\frac{2(n-1)C^{2}}{nr^{2(n-1)}}+\frac{Q}{r^{4q-2}}\\ &Q=\frac{[n(n-1)Q_{1}^{2}]^{q}}{n(4q-n-1)}.\end{split}$ (3.5)
@@
It should be highlighted that within this framework, the variable $(m)$ corresponds to the black hole’s mass. The symbols $(C)$ and $(Q_{1})$ are indicative of the charges associated with the Maxwell field and the Yang-Mills field, respectively. When considering the cosmological constant as a thermodynamic variable in an extended phase space, it is interpreted as the thermodynamic pressure $(P)$, defined by the relation $(P=-\frac{\Lambda}{8\pi})$. Under these stipulations, the expressions for the Hawking temperature, the black hole’s mass, and its entropy can be derived as follows[65],
$\begin{split}T=-\frac{C^{2}(n-1)^{2}}{2\pi nr_{+}^{2n-1}}+\frac{2}{3}(n+1)Pr_{+}-\frac{Q(-n+4q-1)}{4\pi r_{+}^{4q-1}}+\frac{n-1}{4\pi r_{+}},\end{split}$ (3.6)
$\begin{split}M=\frac{n\omega_{n}}{48\pi}\bigg{(}\frac{6C^{2}(n-1)}{nr_{+}^{n-1}}+8\pi Pr_{+}^{n+1}+3Qr_{+}^{n-4q+1}+3r_{+}^{n-1}\bigg{)},\end{split}$ (3.7)
@@
$\begin{split}S=\frac{\omega_{n}r_{+}^{n}}{4},\qquad\omega_{n}=\frac{2\pi^{\frac{n+1}{2}}}{\Gamma\left(\frac{n+1}{2}\right)},\end{split}$ (3.8)
where $S$ is entropy. In this context, the term $(r_{+})$ signifies the radius of the black hole’s event horizon, while $(\omega_{n})$ represents the volume of an ( n )-dimensional unit sphere. To determine the critical points, we can utilize the relations $\frac{\partial T}{\partial r_{+}}=\frac{\partial^{2}T}{\partial r_{+}^{2}}=0$. Consequently, the critical temperature and pressure for the AEMPYM black hole are calculated as follows[65],
$T_{c}=\frac{r_{+}^{-2n-4q-1}\bigg{(}(n-1)r^{4q}\big{(}r_{+}^{2n}-2C^{2}(n-1)r_{+}^{2}\big{)}-2qQ(-n+4q-1)r_{+}^{2n+2}\bigg{)}}{2\pi}$
and
$P_{c}=\frac{3\bigg{(}\frac{(n-1)r_{+}^{-2(n+1)}\big{(}nr_{+}^{2n}-2C^{2}(2n^{2}-3n+1)r^{2}\big{)}}{n}+(4q-1)Q(n-4q+1)r_{+}^{-4q}\bigg{)}}{8\pi(n+1)}$